hi i'm having an extreme amount of trouble setting up word problems for linear equations , especially this one .

Your teacher is giving you a test worth 100 points containing 40 questions. There are two-point and four-point questions on the test. How many of each type of question are on the test?

Answer by Alan3354(69443)

hi i'm having an extreme amount of trouble setting up word problems for linear equations , especially this one .
Your teacher is giving you a test worth 100 points containing 40 questions. There are two-point and four-point questions on the test. How many of each type of question are on the test?
-----------------
Word problems are the scourge of the known universe. Maybe humans made a mistake by learning language. Anyway,
to save typing, use t for the two point questions and f for the four point questions.
t+f = 40 (there are 40 questions total)
2t+4f = 100 (2 points for each t, 4 for each f, 100 total)
Now, it's a math problem, 2 equations in 2 unknowns.
----------------
t+f = 40 (no talking)
2t+4f = 100 --> t+2f = 50 (divided by 2)
t+f = 40 (still)
t+2f = 50
---------------- Subtract
0t-f = -10
f = 10 (10 4-point questions)
t = 30 (must be 30, there are 40 total)